# Formation of extremal Reissner-Nordström black holes: insights from numerics

### v1

@@ 4 Discussion
The gluing solutions are strongly dependent on the Ansatz used for the scalar field profile along the null cone. For a given Ansatz, extremality can be reached only if $\mathfrak{e}M$ is sufficiently large, and $\mathfrak{m}/\mathfrak{e}$ sufficiently small. We have obtained lower and upper bounds, respectively, for these parameters, for several Ansätze, up to the $C^{2}$ gluing problem. Our results are summarised in Table 1. When the scalar field is massless, the lower bound on $\mathfrak{e}M$ is determined by the requirement that $\inf r>0$. This is true for all Ansätze we have studied, for any value of the cosmological constant. As shown in Table 1, a positive cosmological constant reduces the lower bound on $\mathfrak{e}M$. A negative cosmological constant instead raises this bound. The bound is well approximated by $\mathfrak{e}Mr_{+}^{2}/QM={\rm constant}$ (this expression is exact for $C^{0}$ gluing). The sign and size of $\Lambda M^{2}$ does not qualitatively change the remaining properties of the gluing solutions compared with the $\Lambda=0$ case. Gluing solutions of different regularity and derived from different Ansätze generally behave in the same (qualitative) way. The exception is when the scalar field is massive. In this case, for $C^{1}$ and higher gluing, different Ansätze and different branches of the same Ansatz may behave in different ways. Overall, there do not appear to be major qualitative differences between $C^{1}$ and $C^{2}$ gluing, apart from the presence of additional solution branches in the latter case.
@@
We expect that the results for $C^{0}$, $C^{1}$ and $C^{2}$ gluing broadly follow through to higher regularity classes too. It is likely that these will involve multiple branches. We expect that these will behave similarly to branches in the lower regularity problems. This motivates the following question: is it only in the Einstein-Maxwell-charged scalar field model that low-regularity gluing qualitatively captures the features of gluing with higher regularity, or is this a generic feature of characteristic gluing? An answer in the affirmative would be very interesting, as it is significantly easier to study the $C^{0}$ and $C^{1}$ cases, even compared to only $C^{2}$. Note that a classical ($C^{2}$) solution to the Einstein equations in the entire spacetime requires $C^{3}$ characteristic data, as the wave equation loses one derivative at the centre $r=0$ [Kehle:2022uvc]. Proving that an analysis of $C^{0}$ gluing, or perhaps more likely $C^{1}$ gluing, is sufficient to obtain a qualitative understanding of the $C^{3}$ problem would therefore be very beneficial.
@@
One goal of this paper was to determine the minimum value of $\mathfrak{e}M$ required to achieve extremality for a given Ansatz (for, say, $\Lambda=\mathfrak{m}=0$). Our numerical methods have shown that, for any given Ansatz, this minimum value is significantly lower than the (large) value required in the proof of [Kehle:2022uvc]. Of course, the most interesting question is what is the minimum value of $\mathfrak{e}M$ when we extremize over all possible Ansätze. In particular, is this minimum value zero, in other words, for any fixed $\mathfrak{e}$ is it possible to form an arbitrarily small extremal black hole in finite time? [^8] We thank Christoph Kehle for raising this question. Or if the minimum value is non-zero then does the answer depend on $k$? This is clearly a more difficult problem than the one we have tackled. But maybe it can be investigated numerically by devising a method to minimize $\mathfrak{e}M$ for an Ansatz depending on a large number of parameters.
A similar question concerns the massive scalar field: if we extremize over all possible Ansätze then what is the minimum value of $\mathfrak{m}/\mathfrak{e}$ required to achieve extremality? The largest values of $\mathfrak{m}/\mathfrak{e}$ for which we have managed to glue to extremal Reissner-Nordström are all well below $1$ so our results are consistent with the proof of [Reall:2024njy] that gravitational collapse to form extremal Reissner-Nordström in finite time is not possible if $\mathfrak{m}/\mathfrak{e}\geq 1$. Perhaps by investigating Ansätze with a large number of parameters it will be possible to explore whether or not this bound is sharp.
All of our Ansätze have the same phase $e^{-iV}$, which was motivated by the proof of [Kehle:2022uvc]. It would be interesting to explore the effects of making different choices for the phase function. The extra freedom in the phase may also help to minimise the value of $\mathfrak{e}M$ and maximise the value of $\mathfrak{m}/\mathfrak{e}$ at which gluing to an extremal black hole is possible.
@@
$C^{1}$ $\min\mathfrak{e}M$ 0 360 1455 39.3 120
$\min\mathfrak{e}M$ $2/9$ 170 686 18.6 56.2
$\max\mathfrak{m}/\mathfrak{e}$ 0 0.03847 0.02193 0.1714
$C^{2}$ $\min\mathfrak{e}M$ 0 2543 504
$\min\mathfrak{e}M$ $2/9$ 1199 238
$\max\mathfrak{m}/\mathfrak{e}$ 0 0.01403
@@
Another objective of this paper was to give concrete examples of gluing solutions produced by the construction of [Kehle:2022uvc]. Having achieved this, the next step would be to give concrete examples of the resulting spacetime solutions. This could be done by using numerical methods to extend our gluing solutions away from the horizon. This would provide insights into the properties of Cauchy data sets whose maximal developments lead to violations of the third law. This might shed light on a question discussed in [Dafermos:2025int], namely whether initial data leading to the formation, in finite time, of an extremal Reissner-Nordström black hole is in any sense “more fined tuned” than initial data leading to the formation, in finite time, of a Reissner-Nordström black hole with some prescribed subextremal charge to mass ratio.
Finally, it is interesting to ask whether the results presented in this paper and in [Kehle:2022uvc] have analogues for black holes in vacuum. Kehle and Unger conjecture that it is indeed possible to form an extremal Kerr black hole in vacuum, in finite time. If true, such a result would give a disproof of the third law that is independent of any choice of matter model. However, there are (at least) two reasons why investigating this using gluing is a much harder problem than the problem of gluing to extremal Reissner-Nordström. First, we no longer have spherical symmetry: in vacuum the free data for the gluing problem is the shear of the null generators of $\mathcal{C}$, and this is now a function of $3$ coordinates $(V,\theta,\phi)$. So any Ansatz-based approach must work with functions of $3$ coordinates. Second, in the Reissner-Nordström case we have a dimensionless parameter $\mathfrak{e}M$ that can be taken large to make gluing easier. In vacuum there is no analogue of this parameter. For this problem, it is likely to be beneficial to search for a low regularity solution first, even $C^{0}$. Indeed, as we observed in our results, many features of the $C^{0}$ solution persist in the higher regularity solutions. Furthermore, since all higher-regularity solutions must also be $C^{0}$ solutions, the absence of the latter would be strong evidence in favour of a third law for vacuum black holes.
@@ Acknowledgements
We are particularly grateful to Christoph Kehle for his insightful discussions and suggestions, and we thank John R. V. Crump and Ryan Unger for helpful discussions. MG is supported by an STFC studentship and a Cambridge Trust Vice-Chancellor’s Award. HSR and JES are supported by STFC grant ST/X000664/1, and JES is also supported by Hughes Hall College.
@@ Appendix A Numerical Implementation
By far the most challenging aspect of the numerics is optimising the parameters $\alpha$ to obtain a candidate solution. We have opted to employ Broyden’s “good” method [Broyden:1965isd] to tackle this problem. This is a quasi-Newton root-finding method which, instead of explicitly computing the Jacobian, estimates it at each step. We initialise the Jacobian by computing it using finite difference methods.
@@ Appendix B Extension of Kehle-Unger with Cosmological Constant
Kehle and Unger’s construction of gluing data extends easily to the case with non-vanishing cosmological constant. First note that Lemmas 4.4 and 4.5 of [Kehle:2022uvc] are in terms of $r_{+}$, so the same argument works with $\Lambda\neq 0$. Lemma 4.6 also involves $\mathfrak{e}$ and $\mathfrak{q}$. Note that for $\Lambda\neq 0$, $Q(1)=\mathfrak{q}Q_{\rm max}\neq\mathfrak{q}M$, but since $Q_{\rm max}\lesssim 1.07M$ for all $\Lambda M^{2}$, the result holds with a small adjustment of constants. The first part of Lemma 4.7 is identical. With $\Lambda\neq 0$ the initialisation of $\partial_{U}r(0)$ is different, but we still obtain $\partial_{U}r(0)<0$ via (15), assuming (in the case of $\Lambda>0$) that the initial sphere is isometric to a sphere in the de Sitter static patch. Moreover, the extra term in (5f) does not change its parity, so Lemma 4.9 is also identical. Thus the only result that changes is Lemma 4.8, which ensures that $\sup\partial_{U}r<0$.
To prove an extension of Lemma 4.8 we proceed in a similar way (see also [Marin:2024rxt]). Since $r\geq r_{+}/2>0$, due to the previous lemmas, it suffices to show that $\sup r\partial_{U}r<0$. First note that
@@
Clearly when $\Lambda<0$ this makes the RHS further negative than when $\Lambda=0$, and so we have $\sup\partial_{U}r<0$ using the argument of [Kehle:2022uvc]. Thus, we need only consider $\Lambda>0$. Choosing a gauge such that $\Omega=1$ on the horizon, we have
@@
Now $Q(V)\leq Q(1)=\mathfrak{q}Q_{\rm max}\leq c_{0}r(V)$, where $c_{0}=c_{0}(\Lambda M^{2})$ is a constant depending on $\Lambda M^{2}$ and we used $r\geq r_{+}/2$ from Lemma 4.4. This is the same step as in [Kehle:2022uvc]. Moreover, $\Lambda M^{2}<2/9$ so $r^{2}\Lambda\leq 2r^{2}/9M^{2}\leq 2r_{+}^{2}/9M^{2}\leq c_{1}$ for some constant $c_{1}$. (One can choose $c_{1}$ to be independent of $\Lambda M^{2}$.) If $1-Q^{2}/r^{2}-\Lambda r^{2}>0$ then we automatically have $\sup\partial_{U}r<0$. So let us consider the opposite case. Since all terms are bounded we have
@@
Estimations from previous lemmas [Kehle:2022uvc] show that $\partial_{V}r(0)\leq c_{3}\beta^{2}r_{+}\leq c_{4}\mathfrak{q}/\mathfrak{e}$ for constants $c_{3}$, $c_{4}=c_{4}(\Lambda M^{2})$. Finally, using (15),

### v2

@@ 4 Discussion
The gluing solutions are strongly dependent on the Ansatz used for the scalar field profile along the null cone. For a given Ansatz, extremality can be reached only if $\mathfrak{e}M$ is sufficiently large, and $\mathfrak{m}/\mathfrak{e}$ sufficiently small. We have obtained lower and upper bounds, respectively, for these parameters, for several Ansätze, up to the $C^{2}$ gluing problem. Our results are summarised in Table 1. When the scalar field is massless, the lower bound on $\mathfrak{e}M$ is determined by the requirement that $\inf r>0$. This is true for all Ansätze we have studied, for any value of the cosmological constant. As shown in Table 1, a positive cosmological constant reduces the lower bound on $\mathfrak{e}M$. A negative cosmological constant instead raises this bound. The bound is well approximated by $\mathfrak{e}Mr_{+}^{2}/QM={\rm constant}$ (this expression is exact for $C^{0}$ gluing). The sign and size of $\Lambda M^{2}$ does not qualitatively change the remaining properties of the gluing solutions compared with the $\Lambda=0$ case. Gluing solutions of different regularity and derived from different Ansätze generally behave in the same (qualitative) way. (In the $C^{2}$ problem, some solution branches behave differently, but these typically cannot form an extremal black hole.) The exception is when the scalar field is massive. In this case, for $C^{1}$ and higher gluing, different Ansätze and different branches of the same Ansatz may behave in different ways. The main difference between $C^{1}$ and $C^{2}$ gluing is the presence of additional branches in the latter case.
@@
We expect that the results for $C^{0}$, $C^{1}$ and $C^{2}$ gluing broadly follow through to higher regularity classes too. It is likely that these will involve multiple branches. We expect that these will behave similarly to branches in the lower regularity problems. This motivates the following question: is it only in the Einstein-Maxwell-charged scalar field model that low-regularity gluing qualitatively captures the features of gluing with higher regularity, or is this a generic feature of characteristic gluing? An answer in the affirmative would be very interesting, as it is significantly easier to study the $C^{0}$ and $C^{1}$ cases, even compared to only $C^{2}$. Note that a classical ($C^{2}$) solution to the Einstein equations in the entire spacetime requires $C^{3}$ characteristic data, as the wave equation loses one derivative at the centre $r=0$ [12]. Proving that an analysis of $C^{0}$ gluing, or perhaps more likely $C^{1}$ gluing, is sufficient to obtain a qualitative understanding of the $C^{3}$ problem would therefore be very beneficial.
@@
One goal of this paper was to determine the minimum value of $\mathfrak{e}M$ required to achieve extremality for a given Ansatz (for, say, $\Lambda=\mathfrak{m}=0$). Our numerical methods have shown that, for any given Ansatz, this minimum value is significantly lower than the (large) value required in the proof of [12]. Of course, the most interesting question is what is the minimum value of $\mathfrak{e}M$ when we extremize over all possible Ansätze. In particular, is this minimum value zero, in other words, for any fixed $\mathfrak{e}$ is it possible to form an arbitrarily small extremal black hole in finite time? [^10] We thank Christoph Kehle for raising this question. Or if the minimum value is non-zero then does the answer depend on $k$? This is clearly a more difficult problem than the one we have tackled. But maybe it can be investigated numerically by devising a method to minimize $\mathfrak{e}M$ for an Ansatz depending on a large number of parameters.
A similar question concerns the massive scalar field: if we extremize over all possible Ansätze then what is the minimum value of $\mathfrak{m}/\mathfrak{e}$ required to achieve extremality? The largest values of $\mathfrak{m}/\mathfrak{e}$ for which we have managed to glue to extremal Reissner-Nordström are all well below $1$ so our results are consistent with the proof of [17] that gravitational collapse to form extremal Reissner-Nordström in finite time is not possible if $\mathfrak{m}/\mathfrak{e}\geq 1$. Perhaps by investigating Ansätze with a large number of parameters it will be possible to explore whether or not this bound is sharp.
All of our Ansätze have the same phase $e^{-iV}$, which was motivated by the proof of [12]. It would be interesting to explore the effects of making different choices for the phase function. The extra freedom in the phase may also help to minimise the value of $\mathfrak{e}M$ and maximise the value of $\mathfrak{m}/\mathfrak{e}$ at which gluing to an extremal black hole is possible.
@@
$C^{1}$ $\min\mathfrak{e}M$ 0 720 2909 78.5 239
$\min\mathfrak{e}M$ $2/9$ 339 1372 37.2 112
$\max\mathfrak{m}/\mathfrak{e}$ 0 0.02720 0.01550 0.1103
$C^{2}$ $\min\mathfrak{e}M$ 0 4417
$\min\mathfrak{e}M$ $2/9$ 2076
$\max\mathfrak{m}/\mathfrak{e}$ 0 0.00992
@@
Another objective of this paper was to give concrete examples of gluing solutions produced by the construction of [12]. Having achieved this, the next step would be to give concrete examples of the resulting spacetime solutions. This could be done by using numerical methods to extend our gluing solutions away from the horizon. This would provide insights into the properties of Cauchy data sets whose maximal developments lead to violations of the third law. This might shed light on a question discussed in [8], namely whether initial data leading to the formation, in finite time, of an extremal Reissner-Nordström black hole is in any sense “more fined tuned” than initial data leading to the formation, in finite time, of a Reissner-Nordström black hole with some prescribed subextremal charge to mass ratio.
Finally, it is interesting to ask whether the results presented in this paper and in [12] have analogues for black holes in vacuum. Kehle and Unger conjecture that it is indeed possible to form an extremal Kerr black hole in vacuum, in finite time. If true, such a result would give a disproof of the third law that is independent of any choice of matter model. However, there are (at least) two reasons why investigating this using gluing is a much harder problem than the problem of gluing to extremal Reissner-Nordström. First, we no longer have spherical symmetry: in vacuum the free data for the gluing problem is the shear of the null generators of $\mathcal{C}$, and this is now a function of $3$ coordinates $(V,\theta,\phi)$. So any Ansatz-based approach must work with functions of $3$ coordinates. Second, in the Reissner-Nordström case we have a dimensionless parameter $\mathfrak{e}M$ that can be taken large to make gluing easier. In vacuum there is no analogue of this parameter. For this problem, it is likely to be beneficial to search for a low regularity solution first, even $C^{0}$. Indeed, as we observed in our results, many features of the $C^{0}$ solution persist in the higher regularity solutions. Furthermore, since all higher-regularity solutions must also be $C^{0}$ solutions, the absence of the latter would be strong evidence in favour of a third law for vacuum black holes.
@@ Acknowledgements
We are particularly grateful to Christoph Kehle for his insightful discussions and suggestions, and to Carsten Gundlach for noticing errors in an earlier version of this paper. We also thank John R. V. Crump and Ryan Unger for helpful discussions. MG is supported by an STFC studentship and a Cambridge Trust Vice-Chancellor’s Award. HSR and JES are supported by STFC grant ST/X000664/1, and JES is also supported by Hughes Hall College.
@@ Appendix A Numerical Implementation
By far the most challenging aspect of the numerics is optimising the parameters $\alpha$ to obtain a candidate solution. We have opted to employ Broyden’s “good” method [5] to tackle this problem. This is a quasi-Newton root-finding method which, instead of explicitly computing the Jacobian, estimates it at each step. We initialise the Jacobian by computing it using finite difference methods.
@@ Appendix B Extension of Kehle-Unger with Cosmological Constant
Kehle and Unger’s construction of gluing data extends easily to the case with non-vanishing cosmological constant. First note that Lemmas 4.4 and 4.5 of [12] are in terms of $r_{+}$, so the same argument works with $\Lambda\neq 0$. Lemma 4.6 also involves $\mathfrak{e}$ and $\mathfrak{q}$. Note that for $\Lambda\neq 0$, $Q(1)=\mathfrak{q}Q_{\rm max}\neq\mathfrak{q}M$, but since $Q_{\rm max}\lesssim 1.07M$ for all $\Lambda M^{2}$, the result holds with a small adjustment of constants. The first part of Lemma 4.7 is identical. With $\Lambda\neq 0$ the initialisation of $\partial_{U}r(0)$ is different, but we still obtain $\partial_{U}r(0)<0$ via (15), assuming (in the case of $\Lambda>0$) that the initial sphere is isometric to a sphere in the de Sitter static patch. Moreover, the extra term in (5f) does not change its parity, so Lemma 4.9 is also identical. Thus the only result that changes is Lemma 4.8, which ensures that $\sup\partial_{U}r<0$.
To prove an extension of Lemma 4.8 we proceed in a similar way (see also [15]). Since $r\geq r_{+}/2>0$, due to the previous lemmas, it suffices to show that $\sup r\partial_{U}r<0$. First note that
@@
Clearly when $\Lambda<0$ this makes the RHS further negative than when $\Lambda=0$, and so we have $\sup\partial_{U}r<0$ using the argument of [12]. Thus, we need only consider $\Lambda>0$. Choosing a gauge such that $\Omega=1$ on the horizon, we have
@@
Now $Q(V)\leq Q(1)=\mathfrak{q}Q_{\rm max}\leq c_{0}r(V)$, where $c_{0}=c_{0}(\Lambda M^{2})$ is a constant depending on $\Lambda M^{2}$ and we used $r\geq r_{+}/2$ from Lemma 4.4. This is the same step as in [12]. Moreover, $\Lambda M^{2}<2/9$ so $r^{2}\Lambda\leq 2r^{2}/9M^{2}\leq 2r_{+}^{2}/9M^{2}\leq c_{1}$ for some constant $c_{1}$. (One can choose $c_{1}$ to be independent of $\Lambda M^{2}$.) If $1-Q^{2}/r^{2}-\Lambda r^{2}>0$ then we automatically have $\sup\partial_{U}r<0$. So let us consider the opposite case. Since all terms are bounded we have
@@
Estimations from previous lemmas [12] show that $\partial_{V}r(0)\leq c_{3}\beta^{2}r_{+}\leq c_{4}\mathfrak{q}/\mathfrak{e}$ for constants $c_{3}$, $c_{4}=c_{4}(\Lambda M^{2})$. Finally, using (15),